OpenStudy (anonymous):

true or false? 11/2x-2/3x=1/6 is x=29

OpenStudy (anonymous):

is it like this? \[11/2x-2/3x=1/6\]

OpenStudy (anonymous):

Yes thats what it loks like

OpenStudy (anonymous):

have you tried plugging in x and then solving the equation to see if both sides equal each other?

OpenStudy (anonymous):

so how would that look then?

OpenStudy (anonymous):

\[\frac{ 11 }{ 2(29) }-\frac{ 2 }{ 3(29) }=\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

do you know how to solve this now?

OpenStudy (texaschic101):

which one is your problem ? (11/2)x - (2/3)x = 1/6 or 11/(2x) - 2/(3x) = 1/6

OpenStudy (anonymous):

not really i got like 9/-1?

OpenStudy (anonymous):